OpenStudy (anonymous):

solve -3x^2+6x+2≥0 can you please show me how to solve it thanks

myininaya (myininaya):

so we have \[3x^2-6x-2 \le 0\] \[\text{ let } f(x)=3x^2-6x-2\] we need to find when f is 0 first \[3x^2-6x-2=0\] we need to solve this equation for x \[x=\frac{-(-6) \pm \sqrt{(-6)^2-4(3)(-2)}}{2(3)}=\frac{6 \pm \sqrt{36+24}}{6}=\frac{6 \pm \sqrt{60}}{6}\] \[x=\frac{6 \pm 2 \sqrt{15}}{6}=\frac{6}{6} \pm \frac{2 \sqrt{15}}{6}=1 \pm \frac{\sqrt{15}}{3}\] so we have  I chose in a number in each of the intervals divided by the zeros of f so do f(3); f(0); f(-1)